OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) = x^2+6x-8 / x-8 Need help on this too :\

OpenStudy (anonymous):

There isn't one

OpenStudy (anonymous):

there is a slant asymptote though

OpenStudy (anonymous):

If the numerator has a higher power then you don't have a horizontal asymptote

OpenStudy (anonymous):

There is no horizontal asymptote.$$\lim_{x\to\infty}\frac{x^2+6x-8}{x-8}=\lim_{x\to\infty}(2x+6)=\infty$$